### v3

# REAL-Prover: Retrieval Augmented Lean Prover for Mathematical Reasoning

Abstract
Nowadays, formal theorem provers have made monumental progress on high-school and competition level mathematics, but few of them generalize to more advanced mathematics. In this paper, we present REAL-Prover, a new open-source stepwise theorem-prover for Lean 4 to push this boundary. This prover, based on our fine-tuned large language model (REAL-Prover-v1) and integrated with a retrieval system (Leansearch-PS), notably boosts performance on solving college-level mathematics problems. To train REAL-Prover-v1, we developed HERALD-AF, a data extraction pipeline that converts natural language math problems into formal statements, and a new open-source Lean 4 interactive environment (Jixia-interactive) to facilitate synthesis data collection. In our experiments, our prover using only supervised fine-tuning achieves competitive results with a 23.7% success rate (Pass@64) on the ProofNet dataset—comparable to state-of-the-art (SOTA) models. To further evaluate our approach, we introduce FATE-M, a new benchmark focused on algebraic problems, where our prover achieves a SOTA success rate of 56.7% (Pass@64).
keywords:
Neural Theorem Proving, Large Language Models, Formal Reasoning, Proof Search
[…]
68T15, 03B35, 68T20
1 Introduction
The formalization of mathematics has long been a great challenge at the intersection of programming and mathematical logic. Over the past decades, theorem provers have made steady progress. In particular, proof assistants have been used to formalize complex theorems such as the Four-Color Theorem [fourcolourthm] and the Feit-Thompson Odd-Order Theorem [oddthm]. These achievements underscore the potential of formal methods to resolve research-level proofs.
Modern Interactive Theorem Provers (ITPs) like Lean [lean] and Coq [coq] have significantly improved usability and automation, lowering the barrier for adoption. Lean, for example, provides a user-friendly environment and a large community-driven library of mathematics (Mathlib) [The_mathlib_Community_2020]. In early 2025, Lean’s Mathlib had grown to encompass more than 210,000 formally proven theorems in various domains. Meanwhile, utilities like LeanSearch [gao2025semanticsearchenginemathlib4] also emerge to help users navigate easily in these numerous domains and facts, saving their efforts on formal proof drafting.
Figure 1: Overview of the REAL-Prover model training pipeline. The figure illustrates two components: (a) the HERALD-AF pipeline, which translates informal mathematical statements into formal ones; (b) the expert iteration pipeline, which iteratively refines the prover model.
Recent advances in large language models (LLMs) have significantly advanced automated theorem proving, successfully generating formal proofs for high-school-level math competitions [polu2020generativelanguagemodelingautomated, jiang2023draftsketchproveguiding, zhao2023decomposingenigmasubgoalbaseddemonstration, wang2023legoproverneuraltheoremproving, xin2024deepseekprover, wu2024internlm2, xin2024deepseekproverv15harnessingproofassistant, xin2025bfsproverscalablebestfirsttree, ren2025deepseekproverv2advancingformalmathematical]. Nonetheless, extending these approaches to undergraduate and graduate-level mathematics remains a significant challenge due to data sparsity, the intricacies of domain-specific reasoning, and the requirement for precise retrieval and application of extensive lemma libraries, where even minor inaccuracies can compromise the correctness of purely neural proof generation.
To address these challenges, we propose REAL-Prover (REtrieval-Augmented Lean Prover), a retrieval-augmented and stepwise framework for automated theorem proving in Lean. Each model invocation is conditioned on the current proof state and enhanced through the retrieval of semantically relevant theorems from Mathlib. The framework comprises three key components. LeanSearch-PS performs semantic premise selection through a two-stage training strategy that improves both retrieval precision and generalization. HERALD-AF extends the HERALD pipeline to systematically translate informal mathematical statements into verified Lean expressions, enabling scalable data generation. Jixia-Interactive provides a Lean 4 interactive environment that support efficient interaction between prover and compiler in stepwise proving. For evaluation, we construct FATE-M (Formal Algebra Theorem Evaluation–Medium), a benchmark designed for graduate-level reasoning in algebraic domains including groups, rings, and fields. REAL-Prover achieves a state-of-the-art accuracy of 56.7% on FATE-M and demonstrates competitive performance on miniF2F and ProofNet, confirming its effectiveness in complex formal reasoning.
1.1 Related works
Automated Theorem Proving
Automated Theorem Proving in formal languages, particularly within the framework of ITPs, is a rapidly advancing research area in artificial intelligence. LLM-based provers can be broadly categorized by their proof generation strategies: whole-proof generation that aims to produce complete proof in a single pass, and proof-step generation that iteratively refines proofs through ITPs.
The whole-proof generation paradigm focuses on end-to-end synthesis of formal proof scripts from theorem statements. Early attempts like DeepSeek-Prover-V1 [xin2024deepseek] demonstrated the feasibility of this approach in Lean 4. Subsequent improvements in DeepSeek-Prover-V1.5 [xin2024deepseekproverv15harnessingproofassistant] introduced hybrid techniques combining Reinforcement Learning from Proof Assistant Feedback (RLPAF) with Group Relative Policy Optimization [shao2024deepseekmath], along with a ‘truncate-and-resume’ mechanism that allows error correction through partial stepwise verification. Goedel-Prover [lin2025goedel] addressed data scarcity through auto-formalization, first converting natural language mathematics into formal statements before generating proofs through iterative training. While these approaches benefit from large-scale supervised fine-tuning and reduced ITP interaction, they struggle with error accumulation in long logical chains and implicit state inference.
This limitation has driven the development of interactive proof-step generation approaches that leverage ITPs feedback during the proving process. Systems like BFS-Prover [xin2025bfsproverscalablebestfirsttree] employ expert iteration with Best-First Search, using Lean compiler feedback to guide tactic generation. HunyuanProver [li2024hunyuanprover] enhances search efficiency through specialized critic models that evaluate intermediate proof states. The LeanDojo framework introduces ReProver [yang2023leandojotheoremprovingretrievalaugmented], which adopts a distinct retrieval-based approach using Dense Passage Retriever [karpukhin2020dense] to directly predict relevant premises. LeanAgent [kumarappan2024leanagent] extends this paradigm through lifelong learning, maintaining a dynamic knowledge base to progressively master theorem proving across multiple repositories. These interactive methods typically combine expert iteration, enabling real-time adaptation to proof state changes while managing computational complexity through guided tree search.
Formal Statement Benchmark
A fundamental aspect of developing and evaluating theorem provers is the utilization of formal statement benchmarks. Prominent examples include miniF2F [miniF2F], which offers a curated set of Olympiad-level formalization problems, and ProofNet [azerbayev2023proofnetautoformalizingformallyproving], which targets undergraduate-level mathematical reasoning. These benchmarks serve as standardized tools to assess the capabilities of theorem provers in generating correct and complete proofs. As the proficiency of provers advances, there is an increasing need for more challenging benchmarks, such as graduate-level or research-oriented problems, to rigorously evaluate their reasoning depth, adaptability, and ability to handle complex or previously unseen formalizations.
1.2 Our Contributions
Our main contributions are summarized as follows:
  • Retrieval-Augmented Proof System. We present a unified proof system that integrates data collection, training, and inference. The system includes three main components. HERALD-AF provides a pipeline for translating informal mathematical statements into formal Lean expressions. REAL-Prover implements a stepwise theorem-proving framework with LeanSearch-PS for semantic premise selection and employs an expert-iteration training paradigm. Jixia-Interactive facilitates efficient and stable interactions with the Lean 4 prover during both training and inference. The integration of these components significantly enhances proof performance on graduate-level formal mathematics.
  • State-Tactic Pair Dataset. We release a comprehensive dataset consisting of 55k state–tactic pairs generated through our expert iteration process. Each data instance encapsulates a formal proof state, its surrounding context, and the corresponding tactic sequence, providing high-quality supervision for both proof search optimization and policy learning.
  • Specialized Benchmark for Abstract Algebra. We present FATE-M (Formal Algebra Theorem Evaluation–Medium), a specialized benchmark comprising paired informal-formal problems in graduate-level abstract algebra. FATE-M complements existing benchmarks such as miniF2F [miniF2F] and ProofNet [azerbayev2023proofnetautoformalizingformallyproving], establishing an effective framework for rigorous formal reasoning assessment.
2 Method
To construct a high-performance automated prover, we propose REtrieval-Augmented Lean Prover (REAL-Prover), a stepwise proof system for Lean 4. To supply REAL-Prover with a high-quality training corpus, we first introduce Hierarchy and Retrieval-based Translated Lean Dataset Auto-formalization (HERALD-AF) in Section˜2.1, a pipeline that translates informal mathematical problems into formal Lean 4 statements. HERALD-AF converts informal mathematical problems into formal Lean 4 statements, yielding a large corpus of candidate theorems. Since these statements still lack certified proofs, in Section˜2.2, we build a complementary pipeline—again driven by REAL-Prover—that synthesizes and automatically verifies the required proofs, completing the training loop.
2.1 HERALD-AF
Our formal statement generation framework builds upon the Hierarchy and Retrieval-based Translated Lean Dataset (HERALD) [gao2025heraldnaturallanguageannotated], which establishes a protocol translating informal statements into formal statements. Figure˜1(a) schematizes HERALD-AF pipeline.
Before applying HERALD-AF, we perform a preprocessing step to extract informal statements from a natural language mathematics corpus. This begins with the use of regular expressions to identify theorem blocks and exercises within the text. Subsequently, an LLM is employed to generate and refine the extracted informal statements.
Once the informal statements are obtained, we apply HERALD-AF to translate them into formal statements. This translation pipeline comprises three key stages: Auto-Formalization, Auto-Informalization, and LLM Judgement. In the Auto-Formalization stage, an open-source model Herald-translator [gao2025heraldnaturallanguageannotated] is used to generate multiple candidate formal statements from each informal input. During Auto-Informalization, these formal candidates are translated back into natural language using DeepSeek-V3 [deepseekai2025deepseekv3technicalreport]. Finally, in the LLM Judgement stage, a general LLM evaluates whether each back-translated statement faithfully reflects the original input. Only those formal statements that pass this consistency check are retained.
Figure 2: Overview of REAL-Prover. This figure illustrates three main processes: From (a) to (b), the prover passes several tactics to the Lean 4 interactive environment (Jixia-interactive) and receives corresponding scored state nodes. From (b) to (c), the prover selects a state node based on its score and checks whether the proof is complete. From (c) to (a), the prover using LLM with retrieval information from LeanSearch-PS generate new tactics.
2.2 REAL-Prover
[…]
2.2.1 LeanSearch-PS
In our work, premise selection addresses the challenge of identifying the most relevant theorems based on the current Lean proof state. To enhance this process, we developed LeanSearch-PS, a retrieval system built around a high-performance embedding model. This model projects both Lean proof states and Mathlib theorems into a shared semantic vector space. During proof generation, the current proof state is embedded, and a nearest-neighbor search is performed to retrieve the top-k most relevant theorems. These selected premises are then supplied as context to our language model, helping it produce more effective and accurate proof steps.
To train the embedding model, we construct a dataset of $(s,t_{\text{pos}})$ pairs extracted from Mathlib using the Jixia tool [jixiaGitHub]. Here, $s$ denotes a Lean proof state, and $t_{\text{pos}}$ denotes the theorem successfully applied at that step. The training process adopts a two-stage framework inspired by transfer learning strategies commonly used in text retrieval [SFRAIResearch2024]. In both stages, we employ the InfoNCE loss[oord2018representation] defined as
$L(q,k_{i})=-\log\frac{\exp(q\cdot k_{0}/\tau)}{\sum_{i=0}^{n}\exp(q\cdot k_{i}/\tau)},$ (1)
where $k_{0}$ is the embedding of positive example, $k_{1},\ldots,k_{n}$ are the embedding of negative examples, and $\tau$ is a temperature hyperparameter. In the subsequent description, we will omit explicit mention of the embedding for brevity.
The training process consists of the following two stages.
  • Initial Training: We treat the proof state $s$ as the query $q$ and its corresponding theorem $t_{\text{pos}}$ as the positive example $k_{0}$. The in-batch negatives strategy is used, where other premises in the same batch serve as negative examples. This enables contrastive learning of premise relevance across proof states.
  • Hard Negative Enhanced Training: To refine fine-grained discrimination, we leverage the initially trained model to mine hard negatives—challenging but incorrect premises for each state. This yields triplets $(s,t_{\text{pos}},t_{\text{hard\_neg}})$, where $t_{\text{hard\_neg}}$ denotes a mined hard negative premise. During this stage, we use $s$ as $q$, $t_{\text{pos}}$ as $k_{0}$, and $t_{\text{hard\_neg}}$ as $k_{i}$ $(i=1,\ldots,n)$. The resulting model after this phase constitutes our final embedding model.
2.2.2 Expert Iteration
Another key component of REAL-Prover is the prover model within the tactic generator. To train a strong prover model, we combine expert iteration to collect synthetic data, as shown in Figure˜1(b). We begin by fine-tuning a prover model using open-source datasets. This initial model is then employed within REAL-Prover to solve formal statements produced by HERALD-AF, generating additional high-quality data. The newly collected data is incorporated into the supervised fine-tuning dataset to train an improved model. This refined model is then used to tackle challenging formal statements that the previous model could not solve. We update our model continuously in this iteration process.
3 Experiment Setup
3.1 Implementation Detail
REAL-Prover Training
We perform supervised fine-tuning on the base models Qwen2.5-Math-7B [yang2024qwen2] using a learning rate of $5\times 10^{-5}$ with a cosine decay scheduler and a maximum context length of 8192 tokens. Our training data comprises the following sources, totaling 210,420 state-tactic pairs:
  1. Post-processed formal proofs collected via expert iteration. Each proof is parsed into state-tactic pairs using Jixia, and the corresponding prompts are enriched with natural language descriptions, the proof context, the current proof state, and retrieval information from LeanSearch-PS. The dataset comprises 15,071 algebra problems from undergraduate and graduate mathematics texts, and 19,633 NuminaMath problems [numina_math_datasets], resulting in 25,818 and 30,589 state-tactic pairs, respectively, after post-processing.
  2. Post-processed, human-annotated datasets were constructed following the same procedure as expert-iteration proofs. Based on 196 annotated undergraduate-level algebra problems, the datasets contain 18,669 state-tactic pairs. Annotations were provided by undergraduate and graduate students majoring in mathematics, ensuring high quality.
  3. Formal proofs are extracted from Mathlib and parsed into state-tactic pairs using Jixia. Prompt inputs are further augmented with retrieval information from LeanSearch-PS, while natural language descriptions are omitted. The resulting dataset comprises 92,152 state-tactic pairs.
  4. The augmented Lean-Workbook dataset [ying2024leanworkbooklargescalelean]. We enhance this dataset by adding retrieval information from LeanSearch-PS to each example, while keeping the original structure intact. The resulting dataset comprises 43,192 state-tactic pairs.
LeanSearch-PS Training
We develop two dense retrievers from E5-mistral-7b-instruct [wang2023improving] with Tevatron [Gao2022TevatronAE] training framework. For both initial training and hard negative enhanced training, we use LoRA to accelerate training, and set the learning rate as $2\times 10^{-5}$, the batch size as 128, the number of epoch as 1, and other hyperparameters as their default values. Queries are formal statements with a maximum length of 128 and passages are formal theorems with a maximum length of 256. For the hard negative examples, we first embed all statements and theorems with the initial trained embedding model, and then randomly select one passage from the top 30 to top 100 most similar ones for each query as its hard negative premise.
REAL-Prover Inference
In our experiments, considering the overall computational budget, we set the number of passes and samples per step to 64. We set our LLM’s temperature to 1.5. Additionally, we set the hyperparameter $\alpha$ of the score function to 0.5, consistent with the configuration used in BFS-Prover [xin2025bfsproverscalablebestfirsttree]
3.2 Benchmark
As our prover highlights its capabilities on college-level mathematics, we created our own benchmark, FATE-M, which focuses specifically on algebraic formal statements to test model’s performance on rich algebra structures. In addition, we also follow prior work [xin2024deepseekprover, theoremllama, li2024hunyuanprover, lin2025goedel] to evaluate our prover’s general capabilities using the ProofNet and MiniF2F benchmarks.
ProofNet
ProofNet consists of 185 validation problems and 186 test problems, drawn from widely used undergraduate pure mathematics textbooks. The dataset spans a variety of topics, including real and complex analysis, linear algebra, abstract algebra, and topology.
MiniF2F
MiniF2F includes 244 validation problems and 244 test problems, collected from a diverse set of mathematical sources such as the AIME, AMC, and IMO competitions. This benchmark focuses on Olympiad-level problems covering core areas of elementary mathematics, including algebra, number theory, and mathematical induction.
FATE-M (Formal Algebra Theorem Evaluation-Medium)
FATE-M is a benchmark designed to evaluate theorem-proving capabilities in Lean4 for undergraduate-level abstract algebra. It consists of 141 problems formalized in Lean4, most accompanied by natural language comments in English or Chinese. Focusing on core topics in abstract algebra, FATE-M spans problems of simple to moderate difficulty, all of which rely on definitions and lemmas already formalized in Lean’s Mathlib. To our knowledge, it is the first benchmark specifically targeting this domain, offering a specialized tool for assessing formal reasoning in algebra. Future expansions of the FATE series will include more advanced topics, such as commutative algebra, with increased difficulty.
The benchmark was created by extracting problems from 12 university-level abstract algebra textbooks (for a full list, see Appendix˜B), which were then formalized by students. Each formalized statement was rigorously verified by Mathlib contributors and PhD students in algebra to ensure it accurately represents the original mathematical meaning. Additionally, Lean-checked proofs are provided for all statements to guarantee their correctness.
We present two examples to illustrate the nature of the benchmark:
import Mathlib
example {G : Type*} [Group G] {A B : Subgroup G} :
(∃ C : Subgroup G, C.carrier = (A.carrier B.carrier)) A B B A := by
This is a medium-level problem in the benchmark. This problem translates to the following natural language statement: Let $A$, $B$, and $C$ be subgroups of a group $G$. If $C=A\cup B$, then either $A\subseteq B$ or $B\subseteq A$.
import Mathlib
open Classical
[…]
example {G : Type*} {p : ℕ} [Group G] [Fintype G] ($p$ : p.Prime) (ord : Fintype.card G = p ^ 2) : P : Subgroup G, (P.Normal) (Fintype.card P = p) :=
This problem represents one of the more challenging exercises in the benchmark. The natural language translation precisely matches the Lean code comment. An equivalent formulation would be: Let $p$ be a prime number. Then every group of order $p^{2}$ contains a normal subgroup of order $p$.
4 Quantitative Result
Our main results are on ProofNet and FATE-M, both of which consist of college-level mathematics problems that require the use of numerous theorems. To evaluate the generalization ability of our system, we also test it on MiniF2F, a benchmark in a different domain than ProofNet and FATE-M. The results are shown below.
4.1 ProofNet Result
[…]
4.2 FATE-M Result
Table 1: Comparison with state-of-the-art models (all 7B parameters) on ProofNet.
Prove System Sampling Budget ProofNet Test
Whole-proof systems
Goedel Prover [lin2025goedel] 32 15.6 %
DeepSeek-Prover-V1.5-SFT [xin2024deepseekprover] 128 15.9 %
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 128 18.2 %
Tree search systems
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 1 × 3200 21.6 %
REAL-Prover (ours) 64 × 64 23.7 %
The performance on the FATE-M benchmark is presented in Table˜2. Given the availability of open-source models and pipelines, we evaluate only Goedel-Prover and DeepSeek-Prover by using their official repositories. In this experiment, our prover achieves a state-of-the-art result with a Pass@64 of 56.7%. This demonstrates that our prover exhibits strong capability in solving college-level algebraic problems.
4.3 MiniF2F Result
[…]
Table 2: Comparison with other system (all 7B parameters) on FATE-M.
Prove System Sampling Budget FATE-M Test
Whole-proof systems
Goedel Prover [lin2025goedel] 128 18.7 %
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 128 31.2 %
Tree search systems
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 64 × 64 41.8 %
REAL-Prover (ours) 64 × 64 56.7 %
Table 3: Comparison with other system on MiniF2F.
Prove System Sampling Budget MiniF2F Test
Whole-proof systems
DeepSeek-Prover-V1.5-RL [xin2024deepseekprover] 102400 60.2%
Leanabell-Prover-GD-RL [zhang2025leanabell] 128 61.1%
Goedel-Prover [lin2025goedel] 25600 64.7%
STP [dong2502stp] 25600 67.6%
Tree search systems
LLMStep [welleck2023llmstep] 1×32×100 27.9 %
GPT-f [polu2020generativelanguagemodelingautomated] 64×8×512 36.6 %
Hypertree Proof Search [lample2022hypertree] 64 × 5000 41.0 %
DeepSeek-Prover-V1.5-RL + RMaxTS [xin2024deepseekproverv15harnessingproofassistant] 32 × 6400 63.5 %
InternLM2.5-StepProver-BF [wu2024internlm2] 256 × 32 × 600 65.9 %
HunyuanProver v16 + BFS + DC [li2024hunyuanprover] 600 × 8 × 400 68.4 %
BFS-Prover [xin2025bfsproverscalablebestfirsttree] 2048 × 2 × 600 70.8 %
REAL-Prover (ours) 64 × 64 54.1 %
4.4 Ablation Study
Table 4: Ablation study on ProofNet Test and FATE-M Test. REAL-Prover-v1-NoRet w/o LeanSearch refers to use the REAL-Prover-v1-NoRet model without LeanSearch-PS. REAL-Prover-v1 w/ LeanSearch refers to use the REAL-Prover-v1 model with LeanSearch-PS enabled.
Prove System ProofNet Test FATE-M Test
REAL-Prover-v1-NoRet w/o LeanSearch 22.6% 44.7%
REAL-Prover-v1 w/ LeanSearch 23.7% 56.7%
In this section, we will analysis the effective of LeanSearch-PS. To compare with REAL-Prover-v1, we trained a prover model, REAL-Prover-v1-NoRet, which used the same dataset but without retrieval information. And in REAL-Prover, we do not use LeanSearch-PS for REAL-Prover-v1-NoRet model. We evaluate both models on the ProofNet and FATE-M benchmarks with results presented in Table˜4. The results show that REAL-Prover-v1 achieves a performance improvement over REAL-Prover-v1-NoRet on both benchmarks. This demonstrates that the retrieval system enhances the prover’s performance on college-level mathematics problems.
We present a comparison between proofs with and without LeanSearch-PS. In Figure˜3, the proof assisted by LeanSearch-PS successfully retrieves critical lemmas and completes verification, whereas the proof without LeanSearch-PS fails to verify. For problems that can be resolved both with LeanSearch-PS and without LeanSearch-PS, the one with LeanSearch-PS often leads to more accurate and human-readable proof, we provide 2 examples in Appendix˜E to illustrate this observation.
Figure 3: Proof using LeanSearch-PS and retrieval results. The left shows part of our input prompt; the right displays the proof generated by REAL-Prover.
5 Conclusion and Future Work
In this work, we propose REAL-Prover, a stepwise proof system designed to tackle challenging mathematical problems. To support its development, we introduce the HERALD-AF pipeline, which translates informal mathematical statements into formal Lean 4 representations, enabling the construction of a large-scale formal dataset. We also develop a new interactive Lean 4 environment, Jixia-Interactive, which facilitates both the training of our prover model, REAL-Prover-v1, and efficient formal proof generation. Furthermore, we present LeanSearch-PS, a theorem retrieval system that improves theorem selection and boosts overall proof success on college-level mathematical benchmarks. To comprehensively evaluate the prover’s effectiveness, we propose the FATE-M benchmark, on which REAL-Prover achieves a state-of-the-art success rate of 56.7%.
Despite these promising results, several limitations remain. The current model is trained purely with supervised objectives, and we have not yet incorporated reinforcement learning (RL) due to its computational demands. Nevertheless, prior studies suggest that RL can unlock latent reasoning capabilities in large language models, indicating that RL-based fine-tuning (e.g., GRPO [shao2024deepseekmath]) constitutes a promising avenue for enhancing performance. In addition, REAL-Prover emits only the final Lean tactics, whereas several leading provers first generate a natural-language chain of thought (CoT) and then translate it into formal steps, enabling powerful test-time reasoning amplification. Introducing an informal-reasoning layer would let the model draft, filter, and organize candidate arguments before committing to formal tactics, potentially improving both premise selection and high-level proof planning.
Future work will focus on these two complementary directions: incorporating RL-based fine-tuning to deepen reasoning capabilities and generalization, and integrating CoT-driven informal reasoning to bridge natural and formal mathematical discourse. These extensions are expected to yield a more autonomous, interpretable, and robust formal reasoning system, advancing the broader objective of scalable, human-aligned mathematical intelligence.
Acknowledgements
This work is supported in part by National Key R&D Program of China grant 2024YFA1014000, the New Cornerstone Investigator Program, and Ubiquant.
Appendix A Prompt for REAL-Prover
In Lean, a formal proof is a fully constructed proof term that is type-checked and verified by the kernel. It represents a complete and correct derivation of a proposition. The state after tactics refers to the intermediate proof state during tactic-based proof construction. It includes the list of remaining goals and the local context at that point. Relationship: 1. Tactics are procedural tools used to incrementally construct a formal proof. 2. Each tactic transforms the current proof state by solving or reducing goals. 3. The state after a tactic reflects the goals that still need to be proven after that tactic has been applied. 4. Once all goals are solved, Lean assembles the underlying proof terms generated by the tactics into a complete formal proof. 5. This final term is then type-checked by the kernel to ensure correctness. In essence, the state after tactics shows where you are in the process of building a formal proof — it’s a snapshot of what’s left to do before the proof is complete. Here is the FORMAL PROOF before the current state: import Mathlib open BigOperators open Real Nat Topology theorem parallelogram_midpoints_coincide_extracted {A B C D : ℂ} : (A + B + C + D) / 4 = (A + C) / 2 (B + D) / 2 = (A + C) / 2 B.re + D.re = 2 * ((A + C) / 2).re := by simp [add_comm, add_left_comm, add_assoc] simp [div_eq_mul_inv, mul_add, mul_comm, mul_left_comm] intro h h Here is the current STATE: A : B : C : D : h : (A + (B + (C + D))) * 4⁻¹ = (A + C) * 2⁻¹ h : B + D = A + C B.re + D.re = A.re + C.re Please generate a tactic in lean4 to solve the state. And here’re some theorems that may be helpful: ID:0 Formal name:Complex.add_re Informal name: Real Part of Sum of Complex Numbers Formal statement: (z w : ℂ), (z + w).re = z.re + w.re ID:5 Formal name: add_add_add_comm Informal name: Commutativity and Associativity of Addition in Additive Commutative Semigroups Formal statement: {G : Type u_3} [inst : AddCommSemigroup G] (a b c d : G), a + b + (c + d) = a + c + (b + d)
Appendix B Source Textbooks for FATE-M
The FATE-M benchmark draws problems from the following abstract algebra textbooks:
  1. 1.
    Pinter, C. C. A Book of Abstract Algebra (2nd ed., 2010)
  2. 2.
    Fraleigh, J. B. A First Course in Abstract Algebra (8th ed., 2021)
  3. 3.
    Anderson, M. & Feil, T. A First Course in Abstract Algebra: Rings, Groups, and Fields (3rd ed., 2015)
  4. 4.
    Rotman, J. J. A First Course in Abstract Algebra with Applications (3rd ed., 2005)
  5. 5.
    Hodge, J. K., Schlicker, S., & Sundstrom, T. Abstract Algebra: An Inquiry-Based Approach (2014)
  6. 6.
    Herstein, I. N. Abstract Algebra (3rd ed., 1996)
  7. 7.
    Grillet, P. A. Abstract Algebra (2nd ed., 2007)
  8. 8.
    Choudhary, P. Abstract Algebra (2008)
  9. 9.
    Saracino, D. Abstract Algebra: A First Course (2nd ed., 2008)
  10. 10.
    Hungerford, T. W. Abstract Algebra: An Introduction (3rd ed., 2014)
  11. 11.
    Dummit, D. S. & Foote, R. M. Abstract Algebra (3rd ed., 2004)
  12. 12.
    Feng, K. & Zhang, P. 300 Problems in Abstract Algebra (2009)
Appendix C Training Details
REAL-Prover training
REAL-Prover training is performed on $\timesA8$00 GPUs (80GB each) for approximately 18 hours, with hyperparameters detailed in Table˜5.
Table 5: REAL-Prover training hyperparameters
Component Setting
Full Fine Tuning Learning rate : $5\times 10^{-5}$; Scheducler: Cosine Decay
Backbone Qwen2.5-Math-7B
Length Prompt Max Length: 8192
Optimizations bf16, flash_attn
Batch Sizes train_batch_size = 2
Training Schedule 3 Epochs
LeanSearch-PS training
LeanSearch-PS training is performed on $\timesL4$0 GPUs (40GB each) for approximately 12 hours with hyperparameters detailed in Table˜6.
Table 6: LeanSearch-PS training hyperparameters
Component Setting
LoRA Fine Tuning Learning rate : $2\times 10^{-5}$
Backbone E5-mistral-7b-instruct
Optimizations bf16, flash_attn
Batch Sizes train_batch_size = 2
Training Schedule 1 Epoch
Appendix D Broader Impacts
This work explores the use of large language models (LLMs) for automated formal theorem proving. Potential positive impacts include:
  1. 1.
    assisting mathematicians in constructing formal proofs more efficiently by automating tedious or routine steps;
  2. 2.
    facilitating the teaching of formal proof assistants such as Lean in undergraduate and graduate curricula;
  3. 3.
    potentially enabling LLMs to autonomously explore new conjectures or discover novel proofs, contributing to mathematical discovery.
However, several risks and limitations should be acknowledged. First, while LLMs may generate correct formal proofs, these proofs can often be opaque or unintuitive, reducing their explanatory value for human users. Second, excessive reliance on automated proof generation may hinder the development of deep mathematical understanding, particularly in educational settings. Third, the reinforcement learning and training process may be susceptible to reward hacking, especially given the complexity and non-determinism of the Lean environment—raising the risk of generating formally valid but semantically meaningless or misleading proofs that may be overlooked by human reviewers.
Careful evaluation, interpretability mechanisms, and human-in-the-loop verification are essential to ensure that such systems are used responsibly and contribute positively to both research and education.
Appendix E Case study
We provide 2 cases in FATE-M to illustrate that even if both REAL-Prover-v1 with LeanSearch-PS and REAL-Prover-v1-NoRet without Leansearch-PS produce the valid results, the one with premise selection often produce faster and readable proofs, as illustrated in Appendix˜E and Appendix˜E.
example {F : Type*} [Field F] [Fintype F] [IsAlgClosed F] : False := by
set e : Fintype F := Fintype.ofFinite F
set L := (AlgebraicClosure F)
have : Infinite L := by exact IsAlgClosed.instInfinite
have := this
have := Fintype.ofFinite F
exact Fintype.false e
(a) Proof with LeanSearch-PS.
example {F : Type*} [Field F] [Fintype F] [IsAlgClosed F] : False := by
have : Fact (Nat.Prime 2) := by norm_num
exact not_finite F
(b) Proof without LeanSearch-PS.
(c) Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsAlgClosed.instInfinite’ from Mathlib, resulting in a more readable proof.
example {G H : Type*} {p : ℕ} [Group G] [Group H] [Fact p.Prime] (gp : IsPGroup p G) (f : G →* H) (sf : Function.Surjective f) : IsPGroup p H := by
exact IsPGroup.of_surjective gp f sf
(d) Proof with LeanSearch-PS.
example {G H : Type*} {p : ℕ} [Group G] [Group H] [Fact p.Prime] (gp : IsPGroup p G) (f : G →* H) (sf : Function.Surjective f) : IsPGroup p H := by
rw [IsPGroup] at gp
intro g
obtain m, hm :=sf g
rcases gp m with km, hkm
use km
rw [← hm, map_pow, hkm, map_one f]
(e) Proof without LeanSearch-PS.
(f) Compare the proofs with and without LeanSearch-PS. The prover in (a) uses the existing instance ‘IsPGroup.of_surjective’ from Mathlib, resulting in a more readable proof.